OpenStudy (anonymous):

Let f(u,v) = u+v+uv2 and u = u(x,y) = ex +y, v = v(x,y) = xy. Compute by finding the formula for the composite function f(g(x,y)),then computing ∂f/∂x and ∂f/∂x directly.

OpenStudy (anonymous):

its actually like this let f(u,v) = u+v+uv^2 and u = u(x,y) = e^x +y, v = v(x,y) = xy.

OpenStudy (anonymous):

can someone just tell me how to find the composite function f(g(x,y))